# Separable models for dynamic signed networks

### v1

@@ 10 Discussion
This paper introduces a novel separable temporal generative framework for modelling the dynamics of signed networks. By distinguishing between interaction effects and conditional sign processes, our approach preserves the flexibility and interpretability of traditional binary exponential random graph models while remaining consistent with the assumptions of structural balance theory (SBT). To empirically test SBT, we specifically incorporate endogenous effects that capture patterns predicted by the theory. Nonetheless, the framework remains fully compatible with the broader class of binary statistics commonly used in separable temporal ERGM (STERGM) processes, including exogenous and lagged covariates. This separation allows for a deeper understanding of how network relations evolve over time, making it suitable for complex systems in which the interplay of both positive and negative relationships is critical to the system’s overall dynamics.
@@
The flexibility and interpretability of our model open up several avenues for future research. One promising direction involves extending our approach to a continuous-time framework, such as those based on longitudinal ERGMs (Koskinen and Lomi, 2013; Koskinen, Caimo and Lomi, 2015), which could significantly enhance the model’s applicability to real-time, dynamic network data. This extension may involve the development of data augmentation strategies in which MCMC algorithms alternate between sampling from the conditional posterior distribution of the model parameters and generating plausible interaction paths that connect consecutive network snapshots. Another direction concerns the adaptation of the methodology to dynamic weighted signed networks, where the transition probabilities of the signed process follow a partially separable temporal model, as recently proposed by Kei, Chen and Padilla (2023). Crucially, the feasibility of these extensions will depend on the development of more efficient and scalable estimation algorithms. As network data grow in complexity and temporal resolution, computational bottlenecks become increasingly significant. Advancing the underlying inferential algorithm – through improved MCMC schemes, variational approximations (Tan and Friel, 2020) – will be essential to ensure the practical usability of these models in applied settings.

### v2

@@ 11 Discussion
This paper introduces a novel separable temporal generative framework for modelling the dynamics of signed networks. By distinguishing between interaction effects and conditional sign processes, our approach preserves the flexibility and interpretability of traditional binary exponential random graph models while remaining consistent with the assumptions of structural balance theory (SBT).
When modelling signed networks, the separable approach used in this paper is particularly advantageous when formation and dissolution processes are plausibly driven by different mechanisms, and when the goal is to rigorously test structural balance configurations while accounting for the potentially strong influence of overall interaction density on the observed signed edges. In such settings, some signed edges may be unobserved or latent, and failing to account for this can bias the results.
To empirically test SBT, we specifically incorporate endogenous effects that capture patterns predicted by the theory. Nonetheless, the framework remains fully compatible with the broader class of binary statistics commonly used in separable temporal ERGM processes, including exogenous and lagged covariates. This separation allows for a deeper understanding of how network relations evolve over time, making it suitable for complex systems in which the interplay of both positive and negative relationships is critical to the system overall dynamics. When the assumptions underlying separability are not met, alternative approaches that relax one of these assumptions can be employed.
@@
The Bayesian framework also enables assessment of potential violations by comparing posterior predictive performance across alternative models, for instance, by examining whether a model that allows conditional dependence between the processes yields a substantially better fit.
@@
One limitation of this empirical analysis is that the signed relationships in the network are derived from a projection of the underlying bipartite network, based on the similarity of interaction patterns within that same structure (Neal et al., 2024). As such, the sign attributed to edges reflects inferred structural similarity rather than directly observed relational content. A more direct approach would involve the use of longitudinal relational data in which tie signs carry an explicit substantive meaning, for example, recorded friendships and antagonistic relationships between individuals.
The flexibility and interpretability of our model open up several avenues for future research. One promising direction involves extending our approach to a continuous-time framework, such as those based on longitudinal ERGMs (Koskinen and Lomi, 2013; Koskinen et al., 2015), which could significantly enhance the model applicability to real-time, dynamic network data. This extension may involve the development of data augmentation strategies in which MCMC algorithms alternate between sampling from the conditional posterior distribution of the model parameters and generating plausible interaction paths that connect consecutive network snapshots. Another direction concerns the adaptation of the methodology to dynamic weighted signed networks, where the transition probabilities of the signed process follow a partially separable temporal model, as recently proposed by Kei et al. (2023). Crucially, the feasibility of these extensions will depend on the development of more efficient and scalable estimation algorithms. As network data grow in complexity and temporal resolution, computational bottlenecks become increasingly significant. Advancing the underlying inferential algorithm, through improved MCMC schemes, variational approximations (Tan and Friel, 2020), will be essential to ensure the practical usability of these models in applied settings.
12 Acknowledgments
The authors thank Marc Schalberger and Cornelius Fritz for sharing an early version of the ergm.sign package, which enabled us to complete the comparative simulation study presented in this paper. We thank the anonymous reviewers for their insightful comments and constructive suggestions, which have substantially improved the clarity and quality of the manuscript.
13 Data availability
The dataset and code used in this paper are available for download at github.com/acaimo/B2Lstergm.